OpenStudy (anonymous):

9. Write an equation of a circle with the given center and radius. center (–7, –6) and radius 2 (1 point)(x + 7)2 + (y + 6)2 = 4 (x – 7)2 + (y – 6)2 = 2 (x + 7)2 + (y + 6)2 = 2 (x – 7)2 + (y – 6)2 = 4

OpenStudy (campbell_st):

the general form is \[(x - h)^2 + (y - k)^2 = r^2\] where (h, k) is the centre and r is the radius in your question centre(-7,-6) so h = -7 and k = -6 and r = 2. Substitute them for the answer.

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2 (x--7)^2+(y-6)^2=2^2 (x+7)^2+(y+6)^2=4 <---- is this my answer????

OpenStudy (campbell_st):

and that would be correct